# Examining the Source of Defects from a Mechanical Perspective for 3D Anomaly Detection

### v1

@@ Abstract
In this paper, we go beyond identifying anomalies only in structural terms and think about better anomaly detection motivated by anomaly causes. Most anomalies are regarded as the result of unpredictable defective forces from internal and external sources, and their opposite forces are sought to correct the anomalies. We introduced a Mechanics Complementary framework for 3D anomaly detection (MC4AD) to generate internal and external Corrective forces for each point. A Diverse Anomaly-Generation (DA-Gen) module is first proposed to simulate various anomalies. Then, we present a Corrective Force Prediction Network (CFP-Net) with complementary representations for point-level representation to simulate the different contributions of internal and external corrective forces. A combined loss was proposed, including a new symmetric loss and an overall loss, to constrain the corrective forces properly. As a highlight, we consider 3D anomaly detection in industry more comprehensively, creating a hierarchical quality control strategy based on a three-way decision and contributing a dataset named Anomaly-IntraVariance with intraclass variance to evaluate the model. On the proposed and existing five datasets, we obtained nine state-of-the-art performers with the minimum parameters and the fastest inference speed. The source is available at this link.
@@ 1 Introductuon
The task of 3D anomaly detection in industrial manufacturing requires identifying anomalies within 3D point clouds, representing structural deviations from the norm LIN2025103139 ; Li_2024_CVPR ; NEURIPS2023_611b896d ; 2025arXiv250205779J . Existing methods are mainly classified into memory bank methods and reconstruction methods liu2023real3d ; cao2025 ; liao2025robustdistributionalignmentindustrial . Memory bank methods store the normal structural features into the memory bank during the training phase and judge the anomalies by comparing the structural features to be tested with the memory bank features during the testing phase BTF ; ismp ; Zhu_2024 ; liu2023real3d . Reconstruction methods train the model to eliminate the abnormal structures and detect the anomalies by comparing the original and reconstructed structures Li_2024_CVPR ; r3dad ; PO3AD . However, as the motivation we illustrated in Section 2, the existing techniques primarily focus on detecting structural features but do not account for the underlying causes of these anomalies. This paper introduces a novel approach, the Mechanics Complementary Framework for 3D Anomaly Detection (MC4AD), which considers anomalies resulting from mechanical forces acting during production, as presented in Figure 2. These forces, including internal and external defective forces, are modeled and counteracted by corrective forces to restore the point cloud to its normal form.
@@
We propose the Diverse Anomaly-Generation (DA-Gen) module, which simulates a range of anomalies by perturbing the surface normal vectors of 3D objects, enabling more realistic defect generation. Additionally, we present the Corrective Force Prediction Network (CFP-Net), which uses complementary representations to predict corrective forces at the point level. By applying a combined loss function consisting of a symmetric and reconstruction loss function, our approach aims to restore the affected regions of the point cloud. The model’s efficiency is further demonstrated through hierarchical quality control (HQC) strategies, designed to simulate real-world, cost-limited industrial environments. Extensive experiments on multiple datasets show that MC4AD achieves state-of-the-art detection and segmentation performance, as shown in Figure 1. We summarize the contributions:
  1. 1.
@@
    We are the first to consider the causes of these anomalies from a mechanical perspective, considering most abnormalities as the action of the resultant force from internal and external defective forces, and correcting the anomalies by simulating the opposite force.
  2. 2.
@@
    We propose a Mechanics Complementary Framework for 3D Anomaly Detection (MC4AD), which first creates more diverse pseudo anomalies with Diverse Anomaly-Generation, and introduces a CTF-Net that captures complementary features to predict internal and external corrective forces. A new combination of losses is proposed to provide symmetry between the internal and external corrective forces and to optimise the overall restorative results.
  3. 3.
@@
    We simulate more realistic environments for industrial 3D Anomaly Detection, propose a hierarchical quality control based on the three-way decision to simulate multiple detections on real assembly lines, and make a dataset named Anomaly-IntraVariance with intra-class variance to evaluate under cost-limited conditions, which is more meaningful in real industry.
  4. 4.
@@
    Extensive experiments demonstrate the effectiveness of our proposed MC4AD. We obtained 9 SOTAs with minimum parameters and fastest inference speed in detection and segmentation tests on the existing and proposed five datasets across 90 classes.
@@ 2 Preliminaries
2.1 Definition Statement
Definition A1. Given a normal sample as 3-manifold $\textbf{M}\in\mathbb{R}^{3}$, which can be discretized as point cloud $P\text{=}\{\textbf{p}_{i=1}^{n}\}\in\mathbb{R}^{n\times 3}$ with each point $\textbf{p}_{i}=(x_{i},y_{i},z_{i})^{\top}$.
Definition A2. Given a resultant defective force $F_{D}(\textbf{p})\in\mathbb{R}^{3}$, damaging normal object M to anomalous object $\textbf{M}^{\prime}$, which can be decompose as a resultant external defective force $F_{E}(\textbf{p})\in\mathbb{R}^{3}$ and a resultant internal defective force $F_{I}\in\mathbb{R}^{3}$. These forces can be defined as
@@ 2.1 Definition Statement
where $f_{D}(\textbf{p})$, $f_{E}(\textbf{p})$, and $f_{I}(\textbf{p})$ represent the resultant defective, component external defective, and component internal defective force, respectively. $S^{+}$ and $S^{-}$ stand for external and internal planes.
Definition A3. Given a resultant corrective force $F_{C}\in\mathbb{R}^{3}$, restoring anomalous object $\textbf{M}^{\prime}$ to normal object M, which can be decompose as a resultant external corrective force $F_{E}^{\prime}\in\mathbb{R}^{3}$ and a resultant internal corrective force $F_{I}^{\prime}\in\mathbb{R}^{3}$. These forces can be defined as
@@
$=-F_{E}(\textbf{p})-F_{I}(\textbf{p})=-F_{D}(\textbf{p})$ $=\int_{A}-f_{E}(\textbf{p})\,dS^{+}+\int_{S^{-}}-f_{I}(\textbf{p})\,d{S^{-}},$
where $f_{D}^{\prime}(\textbf{p})$, $f_{E}^{\prime}(\textbf{p})$, and $f_{I}^{\prime}(\textbf{p})$ represent the component force, component external Corrective force, and component internal Corrective force, respectively.
@@
$=\textbf{M}+\nabla\textbf{M}_{E}+\nabla\textbf{M}_{I}=\textbf{M}+\nabla\textbf{M}=\textbf{M}^{\prime},$
@@
$=\textbf{M}+\nabla\textbf{M}_{E}^{\prime}+\nabla\textbf{M}_{I}^{\prime}=\textbf{M}^{\prime}+\nabla\textbf{M}^{\prime}=\textbf{M}.$
@@
where $\nabla\textbf{M}=-\nabla\textbf{M}^{\prime}$, meaning: any object damaged by a defective force $F_{D}$ can be restored by the opposite resultant Corrective force $F_{C}$, idealy. We show the symbols used in Appenix C.
2.2 Motivation Statement
Deformation defects such as bulges, cracks, and scratches often appear on product surfaces during industrial production. This can be traced back to unexpected mechanical actions during production and transportation. To the best of our knowledge: Existing methods focus solely on structural anomalies without examining the causes of these anomalies, and a thorough physical analysis of the causes may lead to improved anomaly detection.
@@ 2.2 Motivation Statement
Mechanical actions include the external force $F_{E}(\textbf{p})$ and internal force $F_{I}(\textbf{p})$ mentioned in Definition A2, represented by collisions and gas expansion, respectively. The composition of defective forces $F_{D}(\textbf{p})$ of external and internal forces causes defects mentioned in Definition A2. Anomalies arise from an imbalance in one of the internal or external forces. By simulating a composition Corrective force $F_{C}(\textbf{p})$ opposite to the $F_{D}$ mentioned in Definition A3, applying $F_{D}(\textbf{p})$ to the defect, the defect will ideally be restored as mentioned in Definition A4. Based on the mechanical mechanism of defect formation, the conversion of the defective form $\textbf{M}^{\prime}$ to the standard form M is realized by applying the Corrective force $F_{C}(\textbf{p})$, which is inverse to the defective force $F_{D}(\textbf{p})$. We treat each point of the point cloud as a mechanical node subject to the defect force $F_{D}$: the theoretical force vector of a normal node should be a zero vector, while an abnormal node presents a non-zero vector. By predicting the modified force vector $F_{C}(\textbf{p})$ (which is synthesized by the corresponding external force component $F_{E}^{\prime}(\textbf{p})$ and internal force component $F_{I}^{\prime}(\textbf{p})$) at each node, the mechanical influence of $F_{D}(\textbf{p})$ can be reversed to cancel out, thus restoring the normal geometric characteristics of the abnormal node. The motivation is formalized as:
$Damage\ Source:F_{D}(\textbf{p})$ $=\int_{S^{+}}f_{E}(\textbf{p}),d{S^{+}}+\int_{S^{-}}f_{I}(\textbf{p}),dS^{-},$ (4)
$Corrective\ Object:F_{C}(\textbf{p})$ $=\int_{S^{+}}f_{E}^{\prime}(\textbf{p}),d{S^{+}}+\int_{S^{-}}f_{I}^{\prime}(\textbf{p}),dS^{-},$
$if\ \textbf{p}\ is\ normal\ F_{D}(\textbf{p})$ $=\textbf{0}\ else\neq\textbf{0},\ F_{C}(\textbf{p})=-F_{D}(\textbf{p}).$
@@
where $Damage\ Source$ and $Corrective\ Object$ represent the potential damage at each point, and the goal of the model is to restore the damage. We further describe the theoretical implications and explain the connection to 3D anomaly detection in Appendix B.
3 Related Work
2D Anomaly Detection. In 2D anomaly detection (2D-AD), existing approaches mainly follow two paradigms: Generative and discriminative methods Liu_2024 ; 9849507 ; 10443076 . Generative methods that model normal distributions using networks like Generative Adversarial Networks song2021anoseganomalysegmentationnetwork ; liang2022omni , Autoencoders 9681338 ; Liu_2025 or Diffusion models 9857019 ; he2024diffusion to identify deviations, and discriminative methods that leverage supervised learning to train defect classifiers 9879738 ; spade ; efficientad ; Bergmann_2020_CVPR ; RudWeh2022 . There are morphological differences between images and point clouds, and applying 2D-AD techniques directly to 3D presents challenges.
3D Anomaly Detection. 3D anomaly detection (3D-AD) aims to identify and localize anomalous points or regions within 3D point cloud data AST ; shape ; looking3d ; liang2025fencetheoremdualobjectivesemanticstructure . Existing methodologies are broadly categorized into Feature Embedding-based and Reconstruction-based Methods. Feature Embedding-based Methods, including BTF BTF , M3DM m3dm , Reg3D-AD liu2023real3d , Group3AD Zhu_2024 , and ISMP ismp , utilize pre-trained models or mathematical descriptors to extract discriminative features from normal samples. These features are aggregated into a memory bank to establish normative representations, with anomalies detected through statistical deviations between test samples and the stored feature distributions. Reconstruction-based Methods, as exemplified by IMRNet Li_2024_CVPR , R3D-AD r3dad , and PO3AD PO3AD , adopt encoder-decoder architectures to compress point clouds into latent representations and reconstruct their original geometry liu2024splatposerealtimeimagebasedposeagnostic ; Kruse_2024_CVPR . The large language model (LLM) also demonstrates anomaly detection capabilities on the ZERO-SHOT task 2023arXiv231102782C ; Wang_2025_WACV ; pointad ; AdaCLIP ; ENEL ; cheng2024zeroshotpointcloudanomaly . However, existing methods detect anomaly detection only from the morphological structure and never consider the source of the anomaly. This paper rethinks the source of anomalies from a mechanical perspective.
4 Methodology
Figure 2: Overall Architecture of our MC4AD.
@@ 4.1 Diverse Anomaly-Generation
Effective anomaly elimination in 3D models requires training with realistic pseudo-anomalies that closely resemble real defects in morphology and diversity PNAS ; PO3AD . Existing 3D pseudo-anomaly generation methods are limited by their reliance on simplistic geometric patterns and cannot synthesize certain defect types, such as fine scratches. To address this limitation, we introduced Diverse Anomaly-Generation (DA-Gen) to simulate directional defective forces $F_{D}$ by perturbing surface normal vectors with controlled randomness. While prior studies indicate that anomalies primarily propagate along normal directions, our method introduces slight directional deviations from these normals to enhance realism. These deviations are constrained to remain proximate to the original normal orientation. Additionally, parametric stretching is employed to control defect elongation, enabling the synthesis of narrow non-circular anomalies. Given a point cloud $P$, we partition it into $G$ patches $\hat{P}=\{\hat{P}_{i}\}_{i=1}^{G}\in\mathbb{R}^{G\times K\times 3}$, where each patch contains $K$ points sampled via a hybrid strategy combining random selection and $K$-nearest neighbors. The proposed DA-Gen is applied independently to each patch $\hat{P}_{i}$, formulated as:
$F_{D,i}=\left[\beta_{i}\cdot\lambda_{i}\cdot\nu_{i}+(1-\lambda_{i})\cdot\eta_{i}\right]\cdot\gamma_{i}\cdot\frac{\pi_{i}}{\max(\pi_{i})}\cdot\left(1-\sigma_{i}|\pi_{i}|\right).$ (5)

### v2

@@ Abstract
In this paper, we explore a novel approach to 3D anomaly detection (AD) that goes beyond merely identifying anomalies based on structural characteristics. Our primary perspective is that most anomalies arise from unpredictable defective forces originating from both internal and external sources. To address these anomalies, we seek out opposing forces that can help correct them. Therefore, we introduce the Mechanics Complementary Model-based Framework for the 3D-AD task (MC4AD), which generates internal and external corrective forces for each point. We first propose a Diverse Anomaly-Generation (DA-Gen) module designed to simulate various types of anomalies. Next, we present the Corrective Force Prediction Network (CFP-Net), which uses complementary representations for point-level analysis to simulate the different contributions from internal and external corrective forces. To ensure the corrective forces are constrained effectively, we have developed a combined loss function that includes a new symmetric loss and an overall loss. Notably, we implement a Hierarchical Quality Control (HQC) strategy based on a three-way decision process and contribute a dataset titled Anomaly-IntraVariance, which incorporates intraclass variance to evaluate our model. As a result, the proposed MC4AD has been proven effective through theory and experimentation. The experimental results demonstrate that our approach yields nine state-of-the-art performances, achieving optimal results with minimal parameters and the fastest inference speed across five existing datasets, in addition to the proposed Anomaly-IntraVariance dataset. The source is available at this link.
@@ 1 Introductuon
The task of 3D anomaly detection (3D-AD) in industrial manufacturing requires identifying anomalies within 3D point clouds, representing structural deviations from the norm LIN2025103139 ; Li_2024_CVPR ; NEURIPS2023_611b896d ; 2025arXiv250205779J . Existing methods are mainly classified into memory bank methods liu2023real3d ; ismp ; Zhu_2024 and reconstruction methods r3dad ; Li_2024_CVPR ; liao2025robustdistributionalignmentindustrial . Memory bank methods store the normal structural features into the memory bank during the training phase and judge the anomalies by comparing the structural features to be tested with the memory bank features during the testing phase BTF ; ismp ; Zhu_2024 ; liu2023real3d . Reconstruction methods train the model to eliminate the abnormal structures and detect the anomalies by comparing the original and reconstructed structures Li_2024_CVPR ; r3dad ; PO3AD ; cao2025 .
Motivation. Deformation defects such as bulges, cracks, and scratches often occur on product surfaces during industrial production. These defects can be attributed to unexpected mechanical actions during production and transportation. While existing techniques mostly focus on detecting structural features, they do not consider the underlying causes of these anomalies. To address this issue, this paper designs a novel Mechanics Complementary Model (MCM) to simulate the interaction of these mechanical forces, as shown in Section 3. The proposed framework (MC4AD) employs the MCM for the 3D-AD task, which specifically considers anomalies resulting from mechanical forces acting during production. These forces, including internal and external defective forces, are modeled and counteracted by corrective forces. This process aims to restore the point cloud to its normal form, as illustrated in Figure 1.
@@
In MC4AD, we propose the Diverse Anomaly-Generation (DA-Gen) module, which simulates a variety of anomalies by perturbing the surface normal vectors of 3D objects. This approach allows for more realistic defect generation. In addition, we introduce the Corrective Force Prediction Network (CFP-Net), which employs complementary representations to predict corrective forces at the point level. Besides, we apply a combined loss function that includes both symmetric and reconstruction loss functions, aiming to restore the affected regions of the point cloud. The efficiency of our model is further demonstrated through Hierarchical Quality Control (HQC) strategies, which are designed to simulate real-world, cost-limited industrial environments. Extensive experiments conducted on multiple datasets show that our approach, MC4AD, achieves state-of-the-art (SOTA) performance in detection and segmentation, as illustrated in Figure 1.
Overall, the main contributions of this paper are summarized as follows:
@@
  • To the best of our knowledge, we are the first to examine the causes of structural anomalies from a mechanical perspective. We consider it the result of internal and external defective forces, and we aim to correct them by simulating the opposing force.
@@
  • We propose a novel framework (MC4AD) that creates diverse pseudo anomalies with DA-Gen and introduces a CFP-Net that captures complementary features to predict internal and external corrective forces. A new combination of losses is proposed to optimize symmetry between the internal and external corrective forces.
@@
  • We propose an HQC strategy to simulate multiple detections on real assembly lines and construct a new dataset named Anomaly-IntraVariance to evaluate performance under cost-limited conditions, making it more relevant to real-world industry scenarios.
@@
  • Extensive experiments demonstrate the MC4AD’s effectiveness. We achieved nine state-of-the-art results with minimal parameters and the fastest inference speed in detection and segmentation tests across five existing and proposed datasets involving 90 classes.
2 Related Work
2D Anomaly Detection. The existing 2D approaches mainly follow two paradigms: Generative and discriminative methods Liu_2024 ; 9849507 ; 10443076 . Generative methods that model normal distributions using networks like Generative Adversarial Networks song2021anoseganomalysegmentationnetwork ; liang2022omni , Autoencoders 9681338 ; Liu_2025 or Diffusion models 9857019 ; he2024diffusion to identify deviations, and discriminative methods that leverage supervised learning to train defect classifiers 9879738 ; spade ; efficientad ; Bergmann_2020_CVPR ; RudWeh2022 . There are morphological differences between images and point clouds, and applying 2D-AD techniques directly to 3D presents challenges.
3D Anomaly Detection. The 3D-AD methods aim to identify and localize anomalous points or regions within 3D point cloud data AST ; shape ; looking3d ; liang2025fencetheoremdualobjectivesemanticstructure . Existing methodologies are broadly categorized into feature embedding-based and Reconstruction-based Methods. Feature embedding-based methods extract discriminative features from normal samples and create normal distributions. The features to be tested are used to determine anomalies by their distance from the normal distribution BTF ; Zhu_2024 ; m3dm ; liu2023real3d ; ismp . Reconstruction-based methods compress point clouds into latent representations and reconstruct their original feature liu2024splatposerealtimeimagebasedposeagnostic ; Li_2024_CVPR ; r3dad ; PO3AD ; Kruse_2024_CVPR . The large language model (LLM) also demonstrates anomaly detection capabilities on the zero-shot task 2023arXiv231102782C ; Wang_2025_WACV ; pointad ; AdaCLIP ; ENEL ; cheng2024zeroshotpointcloudanomaly . However, existing methods detect anomalies only from the morphological structure and never consider the source of the anomaly. This paper rethinks the source of anomalies from a mechanical perspective.
@@ 3 Preliminaries
Definition A1. Given a normal sample as 3-manifold $\texttt{M}\in\mathbb{R}^{3}$, which can be discretized as point cloud $P\text{=}\{\texttt{p}_{i=1}^{n}\}\in\mathbb{R}^{n\times 3}$ with each point $\texttt{p}_{i}=(x_{i},y_{i},z_{i})^{\top}$.
Definition A2. Given a resultant defective force $F_{D}(\texttt{p})\in\mathbb{R}^{3}$, damaging normal object M to anomalous object $\texttt{M}^{\prime}$, which can be decompose as a resultant external defective force $F_{E}(\texttt{p})\in\mathbb{R}^{3}$ and a resultant internal defective force $F_{I}\in\mathbb{R}^{3}$. These forces can be defined as
@@
where $f_{D}(\texttt{p})$, $f_{E}(\texttt{p})$, and $f_{I}(\texttt{p})$ represent the resultant defective, component external defective, and component internal defective force, respectively. $S^{+}$ and $S^{-}$ stand for external and internal planes.
Definition A3. Given a resultant corrective force $F_{C}\in\mathbb{R}^{3}$, restoring anomalous object $\texttt{M}^{\prime}$ to normal object M, which can be decompose as a resultant external corrective force $F_{E}^{\prime}\in\mathbb{R}^{3}$ and a resultant internal corrective force $F_{I}^{\prime}\in\mathbb{R}^{3}$. These forces can be defined as
@@
$=-F_{E}(\texttt{p})-F_{I}(\texttt{p})=-F_{D}(\texttt{p})$ $=\int_{A}-f_{E}(\texttt{p})\,dS^{+}+\int_{S^{-}}-f_{I}(\texttt{p})\,d{S^{-}},$
where $f_{D}^{\prime}(\texttt{p})$, $f_{E}^{\prime}(\texttt{p})$, and $f_{I}^{\prime}(\texttt{p})$ represent the component force, component external Corrective force, and component internal Corrective force, respectively.
@@
$=\texttt{M}+\nabla\texttt{M}_{E}+\nabla\texttt{M}_{I}=\texttt{M}+\nabla\texttt{M}=\texttt{M}^{\prime},$
@@
$=\texttt{M}+\nabla\texttt{M}_{E}^{\prime}+\nabla\texttt{M}_{I}^{\prime}=\texttt{M}^{\prime}+\nabla\texttt{M}^{\prime}=\texttt{M}.$
@@
where $\nabla\texttt{M}=-\nabla\texttt{M}^{\prime}$, meaning: any object damaged by a defective force $F_{D}$ can be restored by the opposite resultant Corrective force $F_{C}$, idealy. We show the symbols used in Appenix C.
@@
Mechanics Complementary Model. Mechanical actions include the external force $F_{E}(\texttt{p})$ and internal force $F_{I}(\texttt{p})$ mentioned in Def. A2, represented by collisions and gas expansion, respectively. The composition of defective forces $F_{D}(\texttt{p})$ of external and internal forces causes defects mentioned in Def. A2. Anomalies arise from an imbalance in one of the internal or external forces. By simulating a composition Corrective force $F_{C}(\texttt{p})$ opposite to the $F_{D}$ mentioned in Def. A3, applying $F_{D}(\texttt{p})$ to the defect, the defect will ideally be restored as mentioned in Def. A4. Based on the mechanical mechanism of defect formation, the conversion of the defective form $\texttt{M}^{\prime}$ to the standard form M is realized by applying the Corrective force $F_{C}(\texttt{p})$, which is inverse to the defective force $F_{D}(\texttt{p})$. We treat each point of the point cloud as a mechanical node subject to the defect force $F_{D}$: the theoretical force vector of a normal node should be a zero vector, while an abnormal node presents a non-zero vector. By predicting the modified force vector $F_{C}(\texttt{p})$ (which is synthesized by the corresponding external force component $F_{E}^{\prime}(\texttt{p})$ and internal force component $F_{I}^{\prime}(\texttt{p})$) at each node, the mechanical influence of $F_{D}(\texttt{p})$ can be reversed to cancel out, thus restoring the normal geometric characteristics of the abnormal node. The motivation is formalized as:
Damage $\texttt{Damage Source}:F_{D}(\texttt{p})$ $=\int_{S^{+}}f_{E}(\texttt{p}),d{S^{+}}+\int_{S^{-}}f_{I}(\texttt{p}),dS^{-},$ (4)
Corrective $\texttt{Corrective Object}:F_{C}(\texttt{p})$ $=\int_{S^{+}}f_{E}^{\prime}(\texttt{p}),d{S^{+}}+\int_{S^{-}}f_{I}^{\prime}(\texttt{p}),dS^{-},$
ifpis $\texttt{if}\ \texttt{p}\ \texttt{is normal}F_{D}(\texttt{p})$ $=\textbf{0}\ \texttt{else}\neq\textbf{0},\ F_{C}(\texttt{p})=-F_{D}(\texttt{p}).$
@@
where Damage Source and Corrective Object represent the potential damage at each point, and the goal of the model is to restore the damage, respectively. We further describe the theoretical implications and explain the connection to 3D anomaly detection in Appendix B.
4 Method
We propose a novel pipeline for 3D anomaly detection as illustrated in Figure 2. 1) Training phrase: The training point cloud is processed by DA-Gen to generate a pseudo-anomaly point cloud and employ CTF-Net to extract the potential damage forces to generate the corresponding corrective forces, a process constrained by the combined loss. 2) Test phrase: CTF-Net processes the test point cloud to generate corrective forces per point, directly converts force magnitudes into point-level anomaly scores, and takes their maximum as the object anomaly score.
Figure 2: Overview of MC4AD. The proposed MC4AD contains three important parts: 1) Diverse Anomaly-Generation (DA-Gen) transfers training normal samples into pseudo-anomaly point clouds. 2) Corrective Force Prediction Network (CFP-Net) extracts the potential damage forces in the input point cloud and generates the corresponding corrective forces. 3) Combined loss is used to constrain CFP-Net to generate corrective forces that conform to mechanical constraints correctly.
@@ 4.1 Diverse Anomaly-Generation
Effective anomaly elimination in 3D models requires training with realistic pseudo-anomalies that closely resemble real defects in morphology and diversity PNAS ; PO3AD . Existing 3D pseudo-anomaly generation methods are limited by their reliance on simplistic geometric patterns and cannot synthesize certain defect types, such as fine scratches. To address this limitation, we introduced Diverse Anomaly-Generation (DA-Gen) to simulate directional defective forces $F_{D}$ by perturbing surface normal vectors with controlled randomness. While prior studies indicate that anomalies primarily propagate along normal directions, our method introduces slight directional deviations from these normals to enhance realism. These deviations are constrained to remain proximate to the original normal orientation. Additionally, parametric stretching controls defect elongation, synthesizing narrow non-circular anomalies. Given a point cloud $P$, we partition it into $G$ patches $\hat{P}=\{\hat{P}_{i}\}_{i=1}^{G}\in\mathbb{R}^{G\times K\times 3}$, where each patch contains $K$ points sampled via a hybrid strategy combining random selection and $K$-nearest neighbors. The proposed DA-Gen is applied independently to each patch $\hat{P}_{i}$, formulated as:
$F_{D,i}=\left[\beta_{i}\cdot\lambda_{i}\cdot\nu_{i}+(1-\lambda_{i})\cdot\eta_{i}\right]\cdot\gamma_{i}\cdot\frac{\pi_{i}}{\max(\pi_{i})}\cdot\left(1-\sigma_{i}|\pi_{i}|\right),$ (5)